OpenStudy (anonymous):

Solve 6 1/9 - 3 1/3 . Show your work? THANKS<33

OpenStudy (anonymous):

Its easier to change it to an improper fraction so what do the 2 fractions turn to

jhonyy9 (jhonyy9):

1 1/2 = (2*1 +1)/2 = 3/2 for an example do you understand it ?